Let a=i-j,b=j-k,c=k-i and d is a unit vector such that a.d=0=[bcd].Find d.

Grade:12

1 Answers

Arun
5 years ago
 
let d = xi + y j + z k
 
as a.d = 0
 
hence
x – y = 0
also
 
d. (b x c) = 0
 
now b x c = i + j + k = 0
 
hence d. (b x c) = x + y + z = 0
 
now you can solve